4 NH3(g) +50,(9) — 4NO(g) + 6H20(9) how many molecules of H, O can be formed from 3.78 mol O,?
Answer #1

From reaction,

Mol of H2O formed = (6/5)*mol of O2 reacted

= (6/5)*3.78 mol

= 4.536 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 4.536 * 6.022*10^23 molecules

number of molecules = 2.732*10^24 molecules

Answer: 2.73*10^24 molecules